# Dual Dimensions Geometric Representation Learning Based Document Dewarping

### v1

@@ Abstract
Deep learning, Geometric representation learning, Document image rectification.
@@ I Introduction
Recently, representation learning-based rectification techniques have significantly enhanced the visualization and readability of distorted document images. These methods employed convolutional neural networks (CNNs) and transformers to learn 2D deformation field (backward map) that maps the warped image to the scanned image, thereby optimizing the geometric dewarping task [17, 18, 19, 20, 21, 22, 23, 24, 25, 26, 27, 28, 29]. For example, DocUNet [17] utilized UNet networks to regress the 2D backward map. DocTr [21] introduced a transformer-based approach to analyze the dependencies among features extracted from warped images. DocReal [28] combined channel and spatial attention to enhance control point prediction to better capture local deformations. These methods utilized CNNs or attention mechanism to directly regress 2D deformation fields.
@@ III PROPOSED METHOD
HV Fusion Module. This module aims to fuse the two geometric features obtained by the segmentation model. Leveraging the pooling operation’s ability to aggregate local information on the feature map and inspired by the superiority of coordinate attention in image classification [44], we propose using 2D average pooling to capture local information along the $X$ and $Y$ spatial directions. Actually, for feature maps, the so-called $X$ and $Y$ also refer to the two dimensions of width and height. The horizontal $\boldsymbol{F}_{h}$ and vertical features $\boldsymbol{F}_{v}$ of the distorted image obtained by the segmentation model correspond to the pooling in the $X$ and $Y$ directions. The module is shown in Figure  3. For $\boldsymbol{F_{h}}$ and $\boldsymbol{F_{v}}$, $X$ and $Y$ two-dimensional average pooling is used to obtain the mixed pooling features:
@@
After obtaining the mixed attention features, let the features return to different directions of independent line features. In this step, we separate them in the channel dimension. $X/Y$ self attention is used to capture the long-distance dependencies in the $X$ and $Y$ directions from $\boldsymbol{F}_{h}$. The same applies to $\boldsymbol{F}_{v}$. $X/Y$ self attention here refers to the attention relationship between the features of two line feature maps in the same direction ($X$ or $Y$). Formulas are as follows, where $Sig(\cdot)$ represents the $Sigmoid$ activation function, which normalizes feature values to $[0,1]$. Concatenate $\boldsymbol{F}_{h}^{\prime}$ and $\boldsymbol{F}_{v}^{\prime}$ in the channel dimension, and then predict the 2D deformation field $\boldsymbol{\hat{G}}$ through the upsampling module. Here we use the upsampling module of DocTr [21], DocGeoNet [24] and FTDR [27].
@@
In general, the average pooling operation can extract more robust feature representations. The horizontal and vertical characteristics of the lines in the distorted document combined with the coordinate information in the X and Y directions in the feature map help to strengthen the geometric representation of the deformed features. Our experimental results prove the effectiveness of this fusion module.
@@ IV-D Experimental results
Figure 6: Visualization of our method for horizontal and vertical line prediction (middle two columns) and rectification results (last column) on DocUNet [17], DIR300 [24] and DocReal [28] benchmarks.
Result on DocReal Benchmark. Table  IV shows the quantitative evaluation results of distorted image correction on the DocReal benchmark [28]. Our method brings significant improvements in the three metrics of MS-SSIM (3.6%), LD (11.6%) and AD (4.6%). However, our method still needs to be further improved in ED and CER. In the light of our analysis, there are always differences in line heights between English and Chinese documents, especially text lines. This is a limitation of our method. Perhaps adaptive learnable pooling to accommodate line heights is desirable. In order to intuitively observe the output results of the model for distorted document correction, we show qualitative visualization effects in Figure  7. To more precisely examine the line results predicted and the rectified output by our model, we visualize these two lines as depicted in Figure 6 on three benchmarks.
@@
IV-E Ablation studies
To validate the efficacy of our proposed D2Dewarp, we conduct ablation experiments. Firstly, we verify the impact of HV Fusion Module on the corrective effect. As shown in Tables  V, overall, the HV Fusion Module has a positive effect on the final dewarping results, especially in improving document readability (ED and CER). The LD and AD indicators have slight fluctuations, indicating that the HV Fusion Module focuses more on document readability, and more effective fusion modules still need to be studied in the future. In addition, we further experimentally demonstrated the impact of either horizontal or vertical line features on the rectification results as shown in Table  VI. The result shows that the distorted image can be rectified even if one of the line is used, but the complementary features and constraints between the two lines can enhance the perception of deformation characteristics.
Limitations. Our method still has some limitations. As mentioned above, distorted images in real scenes are subject to interference from text backgrounds, irregular charts without clear boundaries, which prevent our method may not effectively enhance performance.

### v2

@@ Abstract
Computer vision, Deep learning, Geometric representation learning, Document image rectification.
@@ I Introduction
Recently, representation learning-based rectification techniques have significantly enhanced the visualization and readability of distorted document images. These methods employed convolutional neural networks (CNNs) and transformers to learn 2D deformation field (backward map) that maps the warped image to the scanned image, thereby optimizing the geometric dewarping task [17, 18, 19, 20, 21, 22, 23, 24, 25, 26, 27, 28, 29, 30, 31, 32]. For example, DocUNet [17] utilized UNet networks to regress the 2D backward map. DocTr [21] introduced a transformer-based approach to analyze the dependencies among features extracted from warped images. DocReal [28] combined channel and spatial attention to enhance control point prediction to better capture local deformations. These methods utilized CNNs or attention mechanism to directly regress 2D deformation fields.
@@ III PROPOSED METHOD
HV Fusion Module. This module aims to fuse the two geometric features obtained by the segmentation model. Leveraging the pooling operation’s ability to aggregate local information on the feature map and inspired by the superiority of coordinate attention in image classification [47], we propose using 2D average pooling to capture local information along the $X$ and $Y$ spatial directions. For feature maps, the so-called $X$ and $Y$ also refer to the two dimensions of width and height. The horizontal $\boldsymbol{F}_{h}$ and vertical features $\boldsymbol{F}_{v}$ of the distorted image obtained by the segmentation model correspond to the pooling in the $X$ and $Y$ directions. The module is shown in Figure  3. For $\boldsymbol{F_{h}}$ and $\boldsymbol{F_{v}}$, $X$ and $Y$ two-dimensional average pooling is used to obtain the mixed pooling features:
@@
After obtaining the mixed attention features $\boldsymbol{F}_{{mix1}}^{\prime}$ and $\boldsymbol{F}_{{mix2}}^{\prime}$, let the features return to different directions of independent line features. In this step, we separate them in the channel dimension. $X/Y$ self attention is used to capture the long-distance dependencies in the $X$ and $Y$ directions from $\boldsymbol{F}_{h}$. The same applies to $\boldsymbol{F}_{v}$. $X/Y$ self attention here refers to the attention relationship between the features of two line feature maps in the same direction ($X$ or $Y$). Formulas are as follows, where $Sig(\cdot)$ represents the $Sigmoid$ activation function, which normalizes feature values to $[0,1]$. Concatenate $\boldsymbol{F}_{h}^{\prime}$ and $\boldsymbol{F}_{v}^{\prime}$ in the channel dimension, and then predict the 2D deformation field $\boldsymbol{\hat{G}}$ through the upsampling module. Here we use the upsampling module of DocTr [21], DocGeoNet [24] and FTDR [27]. The 2D deformation field has two channels, mapping the vertical and horizontal position coordinates between the distorted image pixels and the flat image.
@@
In general, the average pooling operation can extract more robust feature representations. The horizontal and vertical characteristics of the lines in the distorted document combined with the coordinate information in the X and Y directions in the feature map help to strengthen the geometric representation of the deformed features. Our experimental results (Table  V and Table  VI) prove the effectiveness of this fusion module.
@@ IV-D Experimental results
Figure 6: Text-Heavy Dewarped Results Visualization. The input distorted image (first column), horizontal and vertical line prediction (middle two columns) and rectification results (last column) on DocUNet [17], DIR300 [24] and DocReal [28] benchmarks.
Figure 7: Figure-Heavy (text-sparse) Dewarped Results Visualization.
Result on DocReal Benchmark. Table  IV shows the quantitative evaluation results of distorted image correction on the DocReal benchmark [28]. Our method brings significant improvements in the three metrics of MS-SSIM (3.6%), LD (11.6%) and AD (4.6%). However, our method still needs to be further improved in ED and CER. In the light of our analysis, there are always differences in line heights between English and Chinese documents, especially text lines. This is a limitation of our method. Perhaps adaptive learnable pooling to accommodate line heights is desirable. In order to intuitively observe the output results of the model for distorted document correction, we show qualitative visualization effects in Figure  9. To more precisely examine the line results predicted and the rectified output by our model, we visualize these two lines as depicted in Figure 6 on three benchmarks. For text-sparse distorted documents, our method effectively distinguishes document foreground from background using detected document boundaries (Figure  7).
@@
Inference Speed. On the DocUNet benchmark (130 images), our evaluation was conducted on an NVIDIA V100 GPU (32GB) with batch size 16 (24GB memory utilization), completing in 50.13 seconds at an average of 0.39 seconds per image.
@@
IV-E Ablation studies
To validate the efficacy of our proposed D2Dewarp, we conduct ablation experiments. Firstly, we verify the impact of HV Fusion Module on the corrective effect. As shown in Tables  V, overall, the HV Fusion Module has a positive effect on the final dewarping results, especially in improving document readability (ED and CER). The LD and AD indicators have slight fluctuations, indicating that the HV Fusion Module focuses more on document readability, and more effective fusion modules still need to be studied in the future. In addition, we further experimentally demonstrated the impact of either horizontal or vertical line features on the rectification results as shown in Table  VI. In this ablation experiment, whether H-Line or V-Line is used, the same number of self-attention layers is retained to ensure fair comparison. The result shows that the distorted image can be rectified even if one of the line is used, but the complementary features and constraints between the two lines can enhance the perception of deformation characteristics.
To investigate the impact of varying text coverage ratio on image rectification quality, text density was calculated using PaddleOCR on the ground truth images of the DocUNet benchmark (130 images) - as opposed to distorted images - to maximize text detection accuracy. The text coverage ratio was computed as the total area of detected text boxes relative to the entire image area. Images were sorted in ascending order by text coverage and divided into three intervals (from 0.9% to 65%): Low ($\leq 20\%$), Middle (20%-40%), and High ($\geq 40\%$) text density, as detailed in Table  VII. Overall, high text density significantly impacts MS-SSIM, LD, and AD metrics. As text content increases, both MS-SSIM and AD exhibit performance degradation. We attribute this to their computational principles: in high-text-density regions, the discontinuous pixel distribution in text lines (compared to figure-heavy images) increases sensitivity to pixel misalignment. Conversely, LD demonstrates opposite behavior, its performance notably improves with higher text density. This stems from text content’s inherent local characteristics, where abundant local features enhance LD’s feature matching capability.
@@ IV-E Ablation studies
TABLE VII: Impact of Text Coverage Ratio on MS-SSIM, LD, and AD Metrics over the DocUNet [17] benchmark.
[…]
$TCR\leq 20\%$ 0.55 10.57 0.305
$20\%<TCR\leq 40\%$ 0.48 5.97 0.375
$TCR>40\%$ 0.43 5.02 0.418
Limitations. Our method still has some limitations. As mentioned above, distorted images in real scenes are subject to interference from text backgrounds, irregular charts without clear boundaries, which prevent our method may not effectively enhance performance. For this problem, correction performance could potentially be further enhanced by leveraging existing methods to integrate global information such as foreground features or UV maps.